OpenStudy (anonymous):

if JK=7, KH=21, and JL=6, find LI

OpenStudy (anonymous):



OpenStudy (anonymous):

would LI=8

OpenStudy (kc_kennylau):



OpenStudy (kc_kennylau):

Nope

OpenStudy (anonymous):

ok so then how would i solve it

OpenStudy (kc_kennylau):

How did you get 8?

OpenStudy (anonymous):

21/x=7/6 7x=126 x=8

OpenStudy (anonymous):

wait no 18

OpenStudy (kc_kennylau):

Exactly

OpenStudy (anonymous):

lol thanks

OpenStudy (kc_kennylau):

no problem